OpenStudy (anonymous):

Which mixed number falls between three and one-half and five and three-eighths?

OpenStudy (anonymous):

five and five-twelfths three and one-fourth five and one-third three and one-sixth

OpenStudy (anonymous):

plz help me

undeadknight26 (undeadknight26):

im looking...

hero (hero):

There's a way to figure it out without a calculator or picking up a pencil

OpenStudy (anonymous):

how

undeadknight26 (undeadknight26):

I believe its the 2nd one but not sure...

hero (hero):

It's not the second one because 3 and 1/4 is less than 3 and 1/2

OpenStudy (anonymous):

is it c

undeadknight26 (undeadknight26):

So then its the first one since it in between?

hero (hero):

Basically, to figure it out without picking up a pencil, you have to familiarize yourself with categorizing fractions that are closer to zero, closer to 1/2 or closer to 1

hero (hero):

The first choice is correct

undeadknight26 (undeadknight26):

yay!!!

hero (hero):

1/4 is closer to zero 1/6 is even closer to zero 1/2 is half 3/8 and 5/12 are both close to half but 3/8 is closer to half than 5/12 because 3 is closer to 8 than 5 to 12. A fraction such as 7/8 is closer to 1

hero (hero):

1/6 < 2/6 < 3/6 < 4/6 < 5/6

hero (hero):

Remember that order. It will be useful. 0/7< 1/7 < 2/7 < 3/7 < 4/7 < 5/7 < 6/7 < 7/7 0/7 = 0 7/7 = 1

hero (hero):

These are just examples

hero (hero):

What happened to @fireblood72

OpenStudy (anonymous):

can you help me with this Four of Ethan's family members went on a 4 and two-thirds hour road trip. If they all shared the drive time equally, how many hours did each family member drive?

OpenStudy (anonymous):

six-sevenths 1 and two-thirds eleven-twelfths 1 and one-sixth

hero (hero):

Perform this operation: \[\left(4 + \frac{2}{3}\right)\div 4\]

OpenStudy (anonymous):

so 1 1/6?

hero (hero):

Did you do it by hand or by calc?

OpenStudy (anonymous):

calc